OpenStudy (anonymous):

im am sapposed to find the slope with these given points - (-2, -3) and (6, 1) and i got 1/2... is this correct??? im am sapposed to find the slope with these given points - (-2, -3) and (6, 1) and i got 1/2... is this correct??? @Mathematics

OpenStudy (anonymous):

Slope = change in y / change in x = 1 -- 3 / 6 -- 2 = 4/8 = 1/2 So yes.

OpenStudy (anonymous):

cool thanks!(:

OpenStudy (anonymous):

right 8 up 4 slope 4/8=1/2 yup